Using simple interest



Principal Rate Time Interest. Final amount

1200 1% ? 12 1212

100 ? 2 months 5/6 100⅚

? 5% 2 years 5000 55,000


1
Expert's answer
2021-01-14T19:21:42-0500

Principal Rate Time Interest Final

1200 1% 1 year 12 1212

(12/1200 = 0.01 or 1%, so only 1 period passed)

100 5/36% 2 months 5/6 100⅚

(5/6×2/12 = 5/36%)

50,000 5% 2 years 5000 55,000

(55,000 - 5,000 = 50,000).